the coefficient of x^39 in expansion of (x^4-1/x^3)^15 is

Grade:11

1 Answers

Meet
6 years ago
The general term is 15Cr(x^4)^15-r(-x^-3)^r = -15Cr(x)^60-7r so therefore 60-7r=39 from this r=3 so the coefficient of x^39 is - 15C3 = - 455